OpenStudy (anonymous):

Simplify completely: the quantity 14 times x to the 5th power times y to the 4th power plus 21 times x to the third power times y to the 2nd power all over 7 times x to the third power times y

OpenStudy (phi):

\[ \frac{14x^5 y^4 + 21x^3 y^2}{7x^3y} \] ?

OpenStudy (phi):

you can write the problem like this \[ \frac{14x^5 y^4 + 21x^3 y^2}{7x^3y} \\ \frac{14x^5 y^4}{7x^3y}+ \frac{ 21x^3 y^2}{7x^3y} \]

OpenStudy (phi):

can you simplify the numbers part? what is 14/7 ? what is 21/7 ?